OpenStudy (anonymous):

5x + 2 = x - 6 x = 6x - 8 + 2x - 5 = 7 - 2x x =

OpenStudy (anonymous):

dam that's a hard one

OpenStudy (anonymous):

ik :/

OpenStudy (anonymous):

That's true. This one is super hard.

OpenStudy (anonymous):

What don't you understand about these problems?

OpenStudy (anonymous):

5x + 2 = x - 6 -x -x 4x+2=6 -2 -2 4/4x=4/4 x=1

OpenStudy (anonymous):

@KiArUnsYu that is incorrect.

OpenStudy (anonymous):

6x - 8 + 2x - 5 = 7 - 2x combine like terms 8x-13=7-2x +2x =2x 10x-13=7 +13 +13 10/10x=20/10 x=2

OpenStudy (anonymous):

not hard at all.

OpenStudy (anonymous):

6 must remain negative

OpenStudy (anonymous):

for the first one, x=-2

OpenStudy (anonymous):

oh okay so i made a mistake

OpenStudy (anonymous):

The first one is x=-2 The second one is x=2

OpenStudy (anonymous):

dam really?

OpenStudy (anonymous):

yyuuppp